Required information The Briggs and Stratton Commercial Division designs and manufacturers small engines for golf turf maintenance equipment. A robotics-based testing system with support equipment will ensure that their new signature guarantee program entitled "Always Insta-Start" does indeed work for every engine produced.

                      Pull System          Push System

First cost of equipment     $-1,550,000            $-2,250,000

AOC per Year                   $-720,000    $-600,000

Salvage Value                   $110,000               $80,000

Estimated Life                    8 years                8 years

Determine the salvage value for the push system that will make the company indifferent to the two systems. Also, MARR = 8% per year. The salvage value for the push system is determined to be $ in $1000 units.
Answer #1

Present worth (PW) of Pull System ($) = -1,550,000 - 720,000 x P/A(8%, 8) + 110,000 x P/F(8%, 8)

= -1,550,000 - 720,000 x 5.7466** + 110,000 x 0.5403**

= -1,550,000 - 4,137,552 + 59,433

= -5,628,119

If salvage value of Push system be $S, then for equivalence,

PW of Pull System = PW of Push System

-5,628,119 = -2,250,000 - 600,000 x P/A(8%, 8) + S x P/F(8%, 8)

-3,378,119 = -600,000 x 5.7466** + S x 0.5403**

-3,378,119 = -3,447,960 + S x 0.5403

S x 0.5403 = 69,841

S = $129,263 ~ $130,000 (in $'000)